Project

General

Profile

NDOS Commissioning Meeting 2011/11/03

Attendance: Jon, Gavin, Martin, Mark Mathis, Art, Enrique, Gary, Nick, Jan, Minerba, Raphael, Andrew,

Agenda:

  1. Beam and Detector Report
  2. Installation plans
  3. DAQ Update
  4. DCS Update
  5. AOB

Beam and Detector Report

  • Mat:
    • APDs have been stable since the APD redist. finished
    • cooling has been disabled on leaky parylene APD; has not produced real hits in runs (but looks ok in DSO scan). Andrew: have you tried an extended DSO scan?
    • to summarize: 4 parylene, 1 silicone-coated and 2 uncoated APDs are still running fine (cooled)
  • Mat:
    • problems with TDU the past couple of days
    • had to power cycle
    • MIBS parity error was reported today which blocked TDU from sending real signals. Could be problem with either cable or accelerator unit.
    • Qiming has put software fix in, will test on test-stand and then put it into production soon.
  • Andrew:
    • an Ash River production TDU sitting at MINOS building will be removed (tomorrow) so that a PPC board can be put in. Once that's done, the TDU will be put back into the MINOS building.

Installation plans for next week

  • Mat: received new APD mounting hardware, plan to do "dry" install next week. Still waiting for screws; when they come in, Mat will do the dry install
  • Martin: power supply work on the NDOS, but not until Nov. 15. Reps from Wiener will update SW and new power supplies will be put into NDOS for studies. Andrew: have these PSs been tested at UVa? Ralf: yes

DAQ Update

  • Mat: Rick K. is working on DCM firmware for reduced channel count on DCMs. So DCMs will read out only 58 channels. DAQChannelMap will need to be updated.

DCS Update (Martin)

  • no work scheduled for next week

AOB